Suppose that Congress sets the top personal tax rate on interest and dividends at 38% and the top rate on realized capital gains at 17%. The corporate tax rate stays at 21%. Assume capital gains are half of equity income.


a. Compute the difference between the total corporate plus personal taxes paid on debt and the total taxes on equity income if all capital gains are realized immediately. (Do not round intermediate calculations. Round your answer to 4 decimal places.)




b. Compute the difference between the total corporate plus personal taxes paid on debt and the total taxes on equity income if all capital gains are deferred forever. (Do not round intermediate calculations. Round your answer to 4 decimal places.)
Answer #1

Part A

For equity income, with all capital gains realized immediately:

Corporate tax = 0.2100

Personal tax = (0.38*0.5*(1-0.21)) + (0.17*0.5*(1-0.21)) = 0.2173

Total = $0.4273

Part B

Corporate tax = 0.2100

Personal tax = 0.38*0.5*(1-0.21) = 0.1501

Total = $0.3601

Dear Student,

If any doubts then feel free to ask